OpenStudy (anonymous):

help? step by step? (: limit as x approaches 0 of (1-cos(x))/(x^2)

OpenStudy (anonymous):

l'hopital will work nicely first check that if you replace \(x\) by 0 you get \(\frac{0}{0}\) which you do

OpenStudy (anonymous):

then take the derivative top and bottom, bet \[\frac{-\sin(x)}{2x}\] which, if you replace \(x\) by \(0\) still gives \(\frac{0}{0}\) so it is l'hopital again

OpenStudy (anonymous):

this time the derivative gives you \[\frac{-\cos(x)}{2}\] and now replacing \(x\) by \(0\) gives the limit as \(-2\)

OpenStudy (anonymous):

It's actually 1/2 (:

OpenStudy (anonymous):

yeah, that one!

OpenStudy (anonymous):

lol (; I got the point though.